OpenStudy (anonymous):

So I found the mean that was 600 then i did the step where u subtract the numbers by 600 and then I squared the dif and added the numbers and got 1500314 what do i do with this number?

OpenStudy (anonymous):

sorry um im trying to find the standard deviation

OpenStudy (anonymous):

@sue101

OpenStudy (anonymous):

@rizwan_uet

OpenStudy (anonymous):

What you are saying is a little confusing. Why don't you post the question? :)

OpenStudy (anonymous):

Im trying to find the standard deviation and i just need to kno the last step. I found the mean and i subtracted the mean from the numbers then i squared teh difference and added the numbers and got 1500314 now i need to kno wat to do its like n-1=?=? idk what the last step is....

OpenStudy (amistre64):

id suggest working with a "smaller" set .... subtract 100 from all of it

OpenStudy (amistre64):

itll retain the same standard deviation

OpenStudy (anonymous):

idk wat u r saying i just need the last step after adding.

OpenStudy (amistre64):

im saying; make life easier on yourself .... subtract 100 from all the numbers; if your mean is 600, then your modified mean will be 6 2, -2, 3, -14, 1, 10

OpenStudy (anonymous):

wat????

OpenStudy (amistre64):

in fact, this is a good way to find the mean .... since its just going to be 100 more than the smaller set ... im sure im speaking english :/ 102 98 103 86 101 110 -100 -100 -100 -100 -100 -100 ----------------------------------- 2 -2 3 -14 1 10

OpenStudy (amistre64):

2-2+3-14+1+10 = 0 ... that was sweet your original mean should been 100 ....

OpenStudy (anonymous):

no the book told me how to do it... dude please just tell me the last step. its like n-1= something wat is that last step?

OpenStudy (amistre64):

since the mean of this smaller set is 0, all we need to do is subtract 0 from each value ... and square 2^2 = 4 -2^2 = 4 3^2 = 9 -14^2 = 196 1^2 = 1 10^2 = 100 --------------- add: = 314 your set is assumed to be a population, unless it was stated otherwise elsewhere, so divide by n = 6 314/6 = 52.3333 ; sqrt that for the std deviation sqrt(52.3333) = 7.3422

OpenStudy (amistre64):

the numbers YOU found are waaayyyyy off; since the mean is not 600 ....

OpenStudy (texaschic101):

I know what she did....she just added the numbers together....you then need to divide by the total numbers (600/6 = 100). The mean is 100.